# Another look at inference after prediction

### v4

@@ 5.1 Data generation
We next consider two primary scenarios to evaluate the performance of the PB inference methods and to demonstrate the implications of our theoretical results. In both settings, the generative model for the outcome was
$Y=\theta_{X_{1}}X_{1}+\theta_{\text{\tiny{INT}}}X_{1}W_{1}+{\bf{W}}^{\top}\bm{\theta}+\epsilon$
where ${\bf{W}}=(W_{1},\dots,W_{q})^{{\sf\scriptscriptstyle{T}}}$ and the error term $\epsilon$ was simulated independently from a standard normal distribution. We generated $({\bf{W}}^{\top},X_{1})^{{\sf\scriptscriptstyle{T}}}$ from a mean-zero multivariate normal distribution with a compound symmetric covariance structure with unit variance and pairwise correlation of 0.4. Our objective was to perform inference on the marginal effect of $X_{1}$ on $Y$ in a linear model. Specifically, we target the parameter $\beta_{1}^{*}$, defined as
$(\beta_{0}^{*},\beta_{1}^{*})=\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}\left[(Y-(\beta_{0}+\beta_{1}X_{1}))^{2}\right].$
The predictors to train the ML model, ${\bf{Z}}$, were constructed based on ${\bf{W}}$ and $X_{1}$. The scenarios described below were obtained by varying the values of $\theta_{X_{1}}$, $\theta_{\text{\tiny{INT}}}$, and $\bm{\theta}$, as well as the dimension $q$ of ${\bf{W}}$ and the composition of the predictors ${\bf{Z}}$.
Scenario 1 (Strong and Moderate Prediction Models)
We consider two sub-scenarios to illustrate Corollary 1.1, that is, that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the sufficient condition holds. In both sub-scenarios, we let $q=10$, $\theta_{X_{1}}=1$, $\theta_{\text{\tiny{INT}}}=0$, and $\bm{\theta}=(1,1,0.5,0.5,0.25,{\bf{0}}_{1\times 5})^{{\sf\scriptscriptstyle{T}}}$. In Scenario 1a, we obtain a strong prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{1},\dots,W_{10})^{{\sf\scriptscriptstyle{T}}}$ while in Scenario 1b we obtain a weaker prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{5},X_{1}W_{5})^{{\sf\scriptscriptstyle{T}}}$. The $R^{2}$ of $\widehat{Y}$ for $Y$ is 0.9 and 0.6 in Scenarios 1a and 1b, respectively.
Scenario 2 (Weak Prediction Models)
We consider two sub-scenarios to illustrate the results in Corollary 1.1 and Corollary 2.1, which show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with a weak prediction model while $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In both scenarios, we mimic a common scenario where the prediction model does not contain $X_{1}$. In Scenario 2a, we let $q=9$, $\theta_{X_{1}}=1$, $\theta_{\text{\tiny{INT}}}=0$, and $\bm{\theta}=(0.2,0.1\times{\bf{1}}_{1\times 8})^{{\sf\scriptscriptstyle{T}}}$. We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},\dots,W_{9})^{{\sf\scriptscriptstyle{T}}}$. In Scenario 2b, we let $q=3$, $\theta_{X_{1}}=2$, $\theta_{\text{\tiny{INT}}}=0.25$, and $\bm{\theta}=(0.25,0.25,0.25)^{{\sf\scriptscriptstyle{T}}}$. We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},W_{2},W_{3},W_{2}W_{1},W_{3}W_{1},W_{3}W_{2})^{{\sf\scriptscriptstyle{T}}}$. The $R^{2}$ of $\widehat{Y}$ for $Y$ is 0.5 and 0.4 in Scenarios 2a and 2b, respectively.
In all settings, we generated two datasets from the same underlying distribution: (i) a training dataset to train the ML model and (ii) an independent analysis dataset for evaluating the PB inference methods containing a labeled dataset and an unlabeled dataset. We used 10,000 replications of the analysis dataset for each scenario and a single training set as $\hat{f}$ is assumed to be fixed. The size of the training set was 100,000 and the size of the analysis dataset was $n=10,000$. We varied the proportion of labeled data in the analysis dataset by letting $n_{\text{\tiny{lab}}}=300,600,1000$ or $2000$ in each setting.
We evaluated the performance of the PPI and CC methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$. We also included three comparison methods: $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to serve as the baseline, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ to highlight the limitations of naive inference, and a PB inference method from a concurrent pre-print, the PoSt-Prediction Adaptive (PSPA) method.30 The PSPA method is a generalization of the PPI++ method, an extension of the original PPI estimator that adaptively minimizes the trace of the asymptotic variance-covariance of the PPI estimator. The PSPA estimator, denoted as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, is designed to achieve element-wise variance reduction relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ by incorporating the information from $\widehat{Y}$ directly into the estimating equation for $\bm{\beta}^{*}$. We present the percent bias for each estimator as well as the coverage probability (CP) of the 95% confidence intervals (CIs). For the four PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), we also present the relative efficiency (RE) defined as the ratio of the mean squared error (MSE) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ relative to the MSE of the PB inference estimator.
@@ 5.2 Results
Table 1 presents the percent bias of the various estimators across the different scenarios and labeled dataset sizes. With the exception of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, the methods exhibit minimal bias as they are all consistent estimators for the target parameter. The limitations of naive inference are particularly apparent in Scenarios 2a and 2b, which do not include $X_{1}$ in ML model training. $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ exhibits an average percent bias of -47% in Scenario 2a and -64% in Scenario 2b.
Table 2 summarizes the CPs of the 95% CIs for all methods across the various scenarios and labeled dataset sizes. As expected, $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and the four PB inference estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$) maintain coverage near the nominal level in all settings. In contrast, the 95% CIs based on $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ achieve adequate coverage only in Scenario 1a, where a strong prediction model yields negligible bias in estimating $\beta_{1}^{*}$. Coverage drops to 92% in Scenario 1b and falls to 0% in Scenarios 2a and 2b, where bias is substantial. These results underscore the risks of relying on naive inference in practice and highlight the necessity of using reliable PB inference methods.
Figure 2 displays the relative efficiency across the different settings. In Scenario 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ perform similarly, with RE significantly above 1, indicating that the PB inference methods have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Comparing Scenario 1a and 1b, we note a significant decrease in the gains in statistical efficiency of the PB inference methods due to the weaker prediction model used in Scenario 1b. However, all methods are at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, as the sufficient condition in Corollary 1.1 is satisfied. Additionally, as the labeled set size increases and hence $n_{\text{\tiny{lab}}}/n$ increases, the efficiency gain decreases as expected. $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is always at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ in Scenarios 1a and 1b, and more efficient in Scenario 1a when the labeled dataset size is larger due to the difference in the augmentation terms of these two estimators.
Scenario 2 highlights the performance differences between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, which underscore our findings in Corollary 1.1 and Corollary 2.1. In both sub-scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ consistently outperforms both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and is substantially more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2a, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves an average RE of 1.27 across the various sample sizes, indicating a significant gain over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2b, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ again demonstrates superior RE compared to all other methods and remains more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ offers a modest efficiency gain in Scenario 2a and comparable performance to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in Scenario 2b. Both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ are outperformed by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in nearly all scenarios. In Scenario 2a, while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ remains comparable to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, the RE for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is as low as 0.88 with $n_{\text{\tiny{lab}}}=2000$. Both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ demonstrate a substantial loses in efficiency in Scenario 2b, with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ exhibiting a RE of 0.63 with $n_{\text{\tiny{lab}}}=2000$.
Figure 2: Relative efficiency of the PB inference methods compared to baseline estimation with the labeled data ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$). A ratio greater than 1 indicates that the PB inference method has a smaller mean squared error (MSE) than that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
@@ 6 UK Biobank Analysis
We also evaluated the performance of the PB inference methods in an analysis focused on central obesity. Central obesity is excessive accumulation of body fat around the abdominal region and is strongly associated with cardiometabolic diseases such as type 2 diabetes and cardiovascular disease. A major challenge in its study is the limited availability of direct body composition measurements. Android fat mass (AFM), measured via a dual-energy X-ray absorptiometry (DEXA) scan, is a key marker for assessing central obesity in clinical practice. 51 However, due to its high cost, DEXA-derived AFM is usually only measured in a small subset of participants in large-scale studies. For example, in the UK Biobank,52 a prospective cohort study measuring genetic, lifestyle, and health information on approximately 500,000 participants in the UK, DEXA-derived AFM is incompletely ascertained, particularly among populations historically underrepresented in genetic research. 52 For instance, out of $5,397$ South Asian participants in the UK Biobank, only $288$ (5.3%) participants have undergone DEXA scans.
@@
Our goal is to examine associations between AFM and age, sex, and body mass index (BMI) among the South Asian population. Previous research has indicated that this population is at higher risk of central obesity and that it is associated with cardiometabolic diseases at lower BMI thresholds than European populations. 53 Since DEXA-derived AFM is missing for most South Asian participants, it provides a natural opportunity to leverage ML-predicted AFM in downstream analyses. To illustrate the impact of different ML models on the efficiency of the PB inference methods, we trained two random forest models in the White British population ($32,460$ participants) and used them to predict missing AFM values in the South Asian population. We built a “strong” prediction model that incorporated both the covariates of interest (i.e., age, sex, and BMI) as well as additional predictors, including smoking status, alcohol consumption, two bioelectrical impedance measurements of the legs, physical activity scores from the International Physical Activity Questionnaire, and sedentary behavior quantified as total hours spent driving, using a computer, and watching television. We built another “weak” prediction model that included only the additional predictors. The strong prediction model achieved an $R^{2}$ of 0.8 while the weak prediction model had a substantially lower $R^{2}$ of 0.5 in the South Asian subpopulation.
@@
Table 3 presents point estimates, standard errors, and RE, defined as the ratio of the variance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to the variance of the given estimator, across the two prediction model settings for five estimators: the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), and the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$). We do not include naive inference in our analyses due to the limitations demonstrated in our simulation studies. Echoing our theoretical and simulation study results, all methods yield similar point estimates as they are all consistent estimators of the same parameter. In the analysis with the strong prediction model, the predictions provide information on central obesity beyond what is captured by the covariates of interest, satisfying the sufficient condition in Corollary 2.1. Consequently, both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In fact, all four PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$) yield substantially smaller standard errors compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Among the four PB inference methods, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ show comparable performance, with average REs of 1.39, 1.39, and 1.47, respectively. $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ exhibits higher efficiency gains with an average RE of 1.68. The analysis with the weak prediction model highlights the differences in performance between the PPI and CC methods and reinforces the findings in Corollary 2.1 and Corollary 2.2. Here, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ yield larger standard errors than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest, with an average RE of 0.52. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, as well as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, with average RE values of 1.28 and 1.16, respectively. Overall, the CC estimator is the best performing method in terms of efficiency, underscoring its utility in practice.

### v5

@@ 5.1 Data generation
We next evaluate the performance of the CC method and its comparators through extensive simulation studies for both linear and logistic regression. Although the data is generated in a manner similar to prior work in PB inference 22, 51, we consider several additional mechanisms for generating predictions. In all scenarios, we generated four covariates, ${\bf{W}}=(W_{1},W_{2},W_{3},W_{4})$, from a zero-mean multivariate normal distribution with a compound symmetric covariance structure with unit variance and pairwise correlation of 0.4. For linear regression, the true outcome was generated as
$Y=h({\bf{W}})+\epsilon\quad\mbox{where}\quad\epsilon\sim N(0,0.75^{2})$
and $h({\bf{W}})=1+0.1\exp(W_{1})+\sin(W_{2})+W_{3}+W_{3}^{2}+I(W_{4}>0)$. To assess performance of the PB inference methods using different predictions, we generated
$\widehat{Y}=h({\bf{W}})+\tau\quad\mbox{where}\quad\tau\sim N\left(a+b|W_{2}|,\sigma^{2}_{\tau}\right).$
We varied the amount of bias in the prediction by adjusting $a$ and $b$ and the amount of noise by adjusting $\sigma_{\tau}$. The predictions were generated with three types of errors: (i) random error where $a=b=0$ and $\sigma_{\tau}$ is a positive constant, (ii) non-random error where a=2a=-2, $b=0$, and $\sigma_{\tau}$ is a positive constant, and (iii) covariate-dependent error where $a=0$, b=3.5b=-3.5, and $\sigma_{\tau}=c|W_{2}|$ where $c$ is a positive constant. In all three scenarios, $\sigma_{\tau}$ was set to achieve $R^{2}=0.8$ or 0.6 to represent a high and low quality prediction. For logistic regression, we generated the true outcome $Y\sim\text{Bern}\left(\frac{1}{1+\exp[-h({\bf{W}})]}\right).$ The predictions were similarly generated with (i) random error by randomly flipping the true labels with probability $p_{error}$, (ii) non-random error by randomly flipping the labels with $Y=1$ with probability $p_{error}$ and those with $Y=0$ with probability $\frac{p_{error}}{20}$, and (iii) covariate-dependent error by flipping the observations with $Y=1$ with probability $I(W_{4}>0)\times\frac{p_{error}}{4}+I(W_{4}<0)\times 3p_{error}$ and those with $Y=0$ with probability $I(W_{4}>0)\times\frac{p_{error}}{4}+I(W_{4}<0)\times 2p_{error}$. In all three scenarios, $p_{error}$ was set to achieve an overall accuracy of either 0.9 or 0.7 to represent a high and low quality prediction.
For both linear and logistic regression, we focus on making inference for the parameters obtained from regressing $Y$ on ${\bf{X}}=[1,W_{1},W_{3},I(W_{4}>0)]$. The size of the full dataset was $n=10,000$ and the proportion of labeled data was varied by letting $n_{\text{\tiny{lab}}}=300,600,1000$ or $2000$ in each setting. We evaluated the performance of the five aforementioned PB inference methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$. With the exception of the CC method, all estimators were obtained from the ipd R package.52 In the main text, we present the coverage probability (CP) of the 95% confidence intervals (CIs) to evaluate the quality of inferences and the relative efficiency (RE) to assess the improvement over classical inference for each method. The RE is defined as the ratio of the mean squared error (MSE) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ relative to the MSE of the given estimator. In the Appendix, we include results for percent bias and comparisons of the asymptotic standard error (ASE) and empirical standard error (ESE). All of our results are summarized over 2,000 replications.
@@ 5.2 Results
5.2.1 Linear Regression
For brevity, we present results for the lower quality prediction with $R^{2}=0.6$ and the higher quality prediction with $R^{2}=0.8$ in the Appendix. The general patterns for the PB inference methods are consistent across the settings, but with higher gains in efficiency and more pronounced differences across the methods. Figures 2 - 4 display the results for the three types of errors in the predictions. All methods provide coverage close to the nominal level across all scenarios. With respect to RE, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ generally have significantly higher MSE than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ offering some improvement over $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when the proportion of labeled data is larger. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ consistently outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters and in all scenarios. For the prediction with random error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ perform similarly and have REs exceeding 1.5 for all parameters, while $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ is substantially less efficient for the coefficient for $X_{2}$. For the predictions with non-random and covariate-dependent error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ overall outperforms $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$. $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ provides substantial improvement over both methods for the coefficient for $X_{2}$ and over $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ for the coefficients for $X_{1}$ and $X_{3}$. In fact, $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ are both outperformed by $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ for the smaller labeled sample sizes for the coefficient for $X_{2}$.
Figure 2: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for linear regression with a prediction with random error.
Figure 3: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for linear regression with a prediction with non-random error.
Figure 4: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for linear regression with a prediction with covariate-dependent error.
5.2.2 Logistic Regression
We again present results for the lower quality prediction with overall accuracy of 0.7 and the higher quality prediction with 0.9 accuracy in the Appendix. The general patterns are again consistent, but with higher gains in efficiency and more pronounced differences across the methods for the higher quality prediction. Figures 5 - 7 display the results for the three types of errors in the predictions. All methods provide coverage close to the nominal level across all scenarios. With respect to RE, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ underperform relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ while $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ generally have MSE much lower than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in all scenarios. For the predictions with random error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ perform similarly and have REs near 1.2 for all parameters. For the predictions with non-random error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ outperforms $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ for all parameters, offering an additinal 10-20% improvement in efficiency. For the predictions with covariate-dependent error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ performs similarly to $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ for all parameters and is substantially more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ for the coefficient for $X_{3}$.
Figure 5: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for logistic regression with a prediction with random error.
Figure 6: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for logistic regression with a prediction with non-random error.
Figure 7: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for logistic regression with a prediction with covariate-dependent error.
@@ 6 UK Biobank Analysis
We also evaluated the performance of the PB inference methods in an analysis focused on central obesity. Central obesity is excessive accumulation of body fat around the abdominal region and is strongly associated with cardiometabolic diseases such as type 2 diabetes and cardiovascular disease. A major challenge in its study is the limited availability of direct body composition measurements. Android fat mass (AFM), measured via a dual-energy X-ray absorptiometry (DEXA) scan, is a key marker for assessing central obesity in clinical practice. 53 However, due to its high cost, DEXA-derived AFM is usually only measured in a small subset of participants in large-scale studies. For example, in the UK Biobank,54 a prospective cohort study measuring genetic, lifestyle, and health information on approximately 500,000 participants in the UK, DEXA-derived AFM is incompletely ascertained, particularly among populations historically underrepresented in genetic research. 54 For instance, out of $5,397$ South Asian participants in the UK Biobank, only $288$ (5.3%) participants have undergone DEXA scans.
@@
Our goal is to examine associations between AFM and age, sex, and body mass index (BMI) among the South Asian population. Previous research has indicated that this population is at higher risk of central obesity and that it is associated with cardiometabolic diseases at lower BMI thresholds than European populations. 55 Since DEXA-derived AFM is missing for most South Asian participants, it provides a natural opportunity to leverage ML-predicted AFM in downstream analyses. To illustrate the impact of different ML models on the efficiency of the PB inference methods, we trained two random forest models in the White British population ($32,460$ participants) and used them to predict missing AFM values in the South Asian population. We built a “strong” prediction model that incorporated both the covariates of interest (i.e., age, sex, and BMI) as well as additional predictors, including smoking status, alcohol consumption, two bioelectrical impedance measurements of the legs, physical activity scores from the International Physical Activity Questionnaire, and sedentary behavior quantified as total hours spent driving, using a computer, and watching television. We built another “weak” prediction model that included only the additional predictors. The strong prediction model achieved an $R^{2}$ of 0.8 while the weak prediction model had a substantially lower $R^{2}$ of 0.5 in the South Asian subpopulation.
@@
Table 1 presents point estimates, standard errors, and RE, defined as the ratio of the variance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to the variance of the given estimator, across the two prediction model settings for five estimators: the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), and the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$). Echoing our theoretical and simulation study results, all methods yield similar point estimates as they are all consistent estimators of the same parameter. In the analysis with the strong prediction model, the predictions provide information on central obesity beyond what is captured by the covariates of interest, satisfying the sufficient condition in Corollary 2.1. Consequently, both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In fact, all of the PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$) yield substantially smaller standard errors compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Among the PB inference methods, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ show comparable performance, with average REs of 1.39, 1.39, 1.44, and 1.47, respectively. $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ exhibits higher efficiency gains with an average RE of 1.68. The analysis with the weak prediction model highlights the differences in performance between the PPI and CC methods and reinforces the findings in Corollary 2.1 and Corollary 2.2. Here, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ yield larger standard errors than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest, with an average RE of 0.52. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, as well as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, with average RE values of 1.28, 1.29, and 1.16, respectively. Overall, the CC estimator is the best performing method in terms of efficiency, underscoring its utility in practice.